OpenStudy (anonymous):

is this statement true? 2 log 5 x + 1/2 log 5 y - 4 log 5 z = log 5 x^2√(y)/z^4

OpenStudy (linna):

\[\Large 2\log_5(x)+\frac12 \log_5(y)-4\log_5(z)=\log_5\left(\frac{x^2\sqrt y}{z^4}\right)\] is this what you mean?

OpenStudy (anonymous):

no its y-4 but rest is right

OpenStudy (anonymous):

not 4log

OpenStudy (linna):

\[\Large 2\log_5(x)+\frac12 \log_5(y-4)-\log_5(z)=\log_5\left(\frac{x^2\sqrt y}{z^4}\right)\] this?

OpenStudy (anonymous):

yeah but no negative in front of log5 after the (y-4)

OpenStudy (linna):

after (y-4) log... between log and y-4 what sign is there? multiply, plus, minus ?

OpenStudy (linna):

is it plus?